OpenStudy (anonymous):

The y-intercept of the line whose equation is x - y = -1 is -1??

OpenStudy (anonymous):

y =1

OpenStudy (anonymous):

you get y intercept when x=0 0 - y = -1 -y = -1 y = 1 Yes, y = 1!